MAT 141, HONORS CALCULUS, SUNY AT STONY BROOK, FALL 2000

Lecturer: Prof. Christopher Bishop, bishop@math.sunysb.edu, office is 4-112 in Math Building. Office hours are MF 11:40-12:30.

Text: Calculus, 10th edition, Thomas and Finney.

Syllabus: We will cover approximately Chapters 1, 2, 3, 4 and part of Chapter 5 with certain omissions and additions of sections. A more detailed syllabus will be provided.

Classdays. Classes begin on Tuesday, September 5 and end on Friday, December 15.

Holidays: There are no classes on Yom Kipur (Monday, October 9) and Thanksgiving (November 23-25). Wednesday, October 11 will follow a Monday schedule, so class will meet that day.

Lecture times: Monday and Friday, 12:40-2:00 pm.

Lecture location: The lectures will be on Mondays and Fridays, Physics P-118.

Recitation times and locations: There are two recitiations.

• Section 1 meets Tuesday-Thursday at 11:20-12:15 in Humanities 315. The TA is Lee-Peng Teo. Her office number is 2-106 in the Math Building and his email is lpteo@math.sunysb.edu. Her office hours are Tuesdays, 12:40-2:00p.m. in MLC, Wednesdays, 3:00 to 4:00p.m. in the office.

• Section 2 meets Tuesday-Thursday at 5:30-6:20 in Humanities 315. The TA is also Lee-Peng Teo.

Homework. A list of problems from the text and supplementary material will be handed out. The recitation instructor will assign, collect and return homework. He may also give weekly quizzes based on the assigned problems.

Examinations. There will be two midterm tests: Friday, October 13 and Friday, November 17. These will be given during the regular lecture time. The final exam is 11:00-1:30, Friday, December 22. The location will be announced in class during the last week of the semester.

Early Midterm: There will be an evening exam at 8:30 pm on Tuesday, September 14 in a room to be announced. This is part of the placement process and is given to all students in calculus classes. More information and

a practice exam can be found at http://www.math.sunysb.edu/earlyexam

Grades. Grades are computed using: Recitation 10%, Early Midterm 10%, Midterms 20% each, Final 40%.

Disabilities. If you have any condition, such as physical or mental disability, that will make it difficult to complete the course or that will require extra time on examinations, notify your lecturer during the first two weeks of class so that appropriate arrangements can be made.

Math Learning Center: (Physics A-125/127) This is located on the A level of the physics building and provides help with pre-calculus material. The same room also contains the Calculus Resource Room where you can go for help with Calculus classes. It is staffed most days and some evenings; an exact schedule will be posted in the room. Note that this room is closed during final exam week. It also gets crowded before exams, so don’t wait until the last minute to get help there.

Calculus Online: There are many math sites on the internet. For information on the history of Calculus and short biographies of famous mathe-

maticians, try http://www-groups.dcs.st-and.ac.uk/~history and click

on History Topics or the Biography Index. You can find information about math classes at Stony Brook at www.math.sunysb.edu

Drop dates: Last day for Add/Drop is Monday, September 11. Last day to switch to lower numbered math course is Friday, October 27. The last day to withdraw from a course is Friday, November 10.